# Authors’ Instructions

### v3

@@ 6.5 AI for Autonomic Computing: Promising Future Directions
As our systems grow in size and connectedness, AI- and ML-based autonomous computing will become the norm, making the manual handling and customisation of these systems difficult and costly. We anticipate that AI and ML will primarily drive autonomic computing in the coming years, while humans will still have the ability to influence these systems’ actions through carefully designed interfaces. Cyberphysical technologies and digital twins will make quality-assured and mission-critical adaptations imperative [28]. This is why autonomous programmes will be accountable for physical resources, especially processing facility functions.
@@
AI-powered autonomous computing significantly reduces overall maintenance costs. Consequently, spending on maintenance will go down. The number of individuals required to keep the systems running will also decrease. AI-powered autonomous IT systems provide benefits such as reduced time and money spent on installation and upkeep, as well as improved system reliability [33]. The higher-order benefits state that businesses will have an easier time managing their operations when they use IT systems that can adapt to new environments and follow instructions based on company strategy. A further advantage of AI-based autonomous computing is server consolidation, which decreases the need for expensive and labour-intensive human oversight of massive server farms [34]. Employing AI for autonomous computing could make handling systems simpler. Consequently, computer systems will see a substantial improvement. Server load distribution is an additional case that illustrates how applications function. It involves dividing tasks among many servers. Additionally, monitoring energy generation closely is crucial for implementing policies that are both environmentally friendly and economical.
@@
  • Economical: There are benefits to using computer systems rather than physical data centres [35]. Cloud computing makes it easy for businesses to purchase AI technologies, despite the hefty upfront expenditures. AI systems may potentially do data analysis independently of humans.
@@
  • Autonomous: AI-driven cloud computing has the potential to make businesses smarter, more effective, and more insight-driven [36]. Automation of mundane and routine tasks and the ability to undertake data analysis independently of human intervention are two ways in which AI-driven could increase output.
@@
  • Data Management: By combining AI with Google Cloud Stream statistical analysis, it is possible to accomplish real-time customisation, identifying anomalies, and administration scenario forecasting [37].
@@
  • Effective Decisions: The increasing development of cloud applications underscores the importance of intelligence-based data security [38]. Security solutions for networks that use AI make it feasible to trace and analyse network traffic. Systems driven by AI can immediately initiate an alarm response upon detection of an anomaly. This method protects important data.
@@ 7 Conclusions and Summary
The article delves into the relationship between SLA violations and QoS in cloud-based systems, as well as their respective roles. To learn how quality of service needs affect self-management, one must be familiar with the development of autonomous cloud computing in order to ascertain the scheduling efficiency of the provided resources. Ultimately, it identified unanswered questions about autonomous cloud computing and other open research topics, which ML and AI systems could potentially address by anticipating resource needs. Importantly, self-adaptive systems driven by AI offer to be environmentally friendly and economically address evolving needs in an unpredictable environment, rather than merely allocating increasing amounts of resources. Consequently, some industry heavyweights are studying and implementing autonomic computing with cutting-edge AI and ML solutions.

### v4

@@ 7.1 Transformative Effects of AI on Autonomic Cloud Computing
As systems grow in size and connectivity, AI- and ML-based autonomous computing will become the norm, making the manual handling and customisation of these systems difficult and costly. It has been anticipated that AI and ML will primarily drive autonomic computing in the coming years, while humans will still have the ability to influence these systems’ actions through carefully designed interfaces. Cyberphysical technologies and digital twins will make quality-assured and mission-critical adaptations imperative A39 . This is why autonomous programmes will be accountable for physical resources, especially processing facility functions.
@@
AI-powered autonomic computing significantly reduces overall maintenance costs. Consequently, spending on maintenance will go down. The number of individuals required to keep the systems running will also decrease. AI-powered autonomous IT systems provide benefits such as reduced time and money spent on installation and upkeep, as well as improved system reliability A32 . The higher-order benefits state that businesses will have an easier time managing their operations when they use IT systems that can adapt to new environments and follow instructions based on company strategy. A further advantage of AI-based autonomic computing is server consolidation, which decreases the need for expensive and labour-intensive human oversight of massive server farms A33 . Employing AI for autonomic computing could make handling systems simpler. Consequently, computing systems will see a substantial improvement. Server load distribution is an additional case that illustrates how applications function. It involves dividing tasks among many servers. Additionally, monitoring energy generation closely is crucial for implementing policies that are both environmentally friendly and economical.
@@
  • Economical: There are benefits to using computing systems rather than physical data centers A34 . Cloud computing makes it easy for businesses to purchase AI technologies, despite the hefty upfront expenditures. AI systems may potentially do data analysis independently of humans.
@@
  • Autonomous: AI-driven cloud computing has the potential to make businesses smarter, more effective, and more insight-driven A35 . Automation of mundane and routine tasks and the ability to undertake data analysis independently of human intervention are two ways in which AI-driven could increase output.
@@
  • Data Management: By combining AI with Google Cloud Stream statistical analysis, it is possible to accomplish real-time customisation, identifying anomalies, and administration scenario forecasting A36 .
@@
  • Effective Decisions: The increasing development of cloud applications underscores the importance of intelligence-based data security A37 . Security solutions for networks that use AI make it feasible to trace and analyse network traffic. Systems driven by AI can immediately initiate an alarm response upon detection of an anomaly. This method protects important data.
7.2 AI-driven Autonomic Computing: A Conceptual Model
Figure 8 illustrates the conceptual software design, highlighting the connection between AI/ML and the autonomic computing paradigm A26 . To meet demands for a wide range of IoT apps, this model incorporates cutting-edge technology to provide efficient computing solutions.
Figure 8: A Conceptual Model for AI-driven Autonomic Computing A26
1. IoT Applications: By communicating with IT infrastructures, gateways allow users and IoT gadgets to abstract away interactions with edge-based actuators, sensing devices, and effectors A26 . To effectively provide AI and other autonomous services, the system can connect to a variety of devices and applications, including medical care, urban planning, agricultural surveillance, and meteorological monitoring.
2. Cloud Resource Manager: Flexible and resilient resource and task allocation is essential for distributed systems, such as edge platforms of IoT. The resource management module tracks the QoS of each distributed task, the needed resources, limitations (such as location), and the available and allocated resources A26 . In order to produce a workload schedule, the resource manager allocates tasks based on available resources from resource pool and uses AI to estimate their completion time. Using AI and ML approaches, the resource scheduler estimates the system’s resources and efficiently allocates them to complete activities within the given time and cost. In addition, this module integrates provider-supplied data on available and planned resources, along with resource descriptions.
3. Autonomic Model: This next-generation architecture makes use of IBM’s autonomic computing paradigm with AI/ML, which focuses on capabilities like self-optimization, self-protection, self-configuring, and self-healing.
  • Self-* Properties: The goal of self-healing is to rebound from flaws using the latest AI models while maintaining the whole system functioning without disruption by implementing all necessary alterations. By monitoring unusual behavior and acting accordingly, self-protection aims to keep the system safe from malicious intent and ensure its smooth operation. The fundamental objective of self-configuration is to install older or missing components automatically using AI, eliminating the need for human intervention A26 . The self-optimizing part uses dynamic scheduling algorithms to assign tasks and workloads to the most suitable resources. AI-driven adaptive scheduling leverages the autonomous component’s feedback to consistently enhance the system’s efficiency. Data-intensive systems can benefit from AI/ML-based adaptable scheduling due to its adaptability and versatility. In addition, it is possible to automatically assess the effect of various QoS attributes on system efficacy using the latest AI models A49 . Additionally, AI-driven autonomic computing can optimize the distribution of server loads and energy usage, resulting in cost savings and reduced environmental impacts A50 .
  • MAPE-k Loop: Sensors record the current state nodes’ effectiveness in terms of QoS measures. In order to continuously check for performance differences, monitors compare the actual results with those anticipated by AI using data collected from the resource manager node. The monitoring module sends data to the Analyze and Plan component, which processes it and determines how to respond based on potential recommendations from AI models. The executor’s major goal is self-optimization, which allows it to improve QoS and complete tasks within a certain time frame. The effector serves as a conduit for communicating novel rules, guidelines, and warnings to various computing nodes. Along with other QoS needs, the knowledge base stores the primary rules for scheduling resources A26 . AI-powered autonomous agents will assume the role of a system administrator, automating the execution process. Furthermore, AI-driven methods can improve the efficiency of the persistence (knowledge) component of the MAPE-K loop. This is particularly true when it comes to reliably resolving state synchronization in a widely dispersed context.
AI-powered autonomic computing has the capability to significantly improve the overall performance of computing systems.
8 Promising Future Directions
The following are possible future directions that can be addressed by future readers:
8.1 Security and Privacy
Research on IoT, edge computing, and cloud computing has shifted dramatically in recent years toward autonomic computing, all with the goal of better serving consumers A41 . However, this enormous paradigm shift raises numerous challenges in ensuring the privacy and security of data stored on these devices. For autonomic edge computing to work, privacy and security measures must be strong and flexible. This is because autonomic edge computing has some special features, like less delay, geographical distribution, end-device connectivity, high computational power, fluctuations, and more. Due to the heterogeneity of devices and applications, it is also difficult to develop software platforms that are interoperable A42 . The increasing deployment of cloud-based smartphone applications makes intelligence-based autonomous data protection crucial. Security solutions for networks that use AI make it possible to track and analyze network traffic. A self-protection module within autonomic computing can use AI to immediately initiate an alarm response upon detection of an anomaly and protect vital data. Applications’ security is heavily dependent on their edge infrastructure to withstand threats including replay, denial of service attacks, information degradation, and identification revelation. In autonomic computing environments, consumers lack control over their personal information, necessitating the establishment of authorization, encryption, and data aggregation. Data on consumers and vendor partners are just one piece of the massive amount of data needed by AI systems A43 . Knowing who possesses the data is significantly more valuable than having confidential data that cannot be associated with a specific individual. Companies often face challenges with data security and regulatory compliance when using sensitive content. Privacy and data law security must be considered when implementing AI in autonomic computing A44 . Given the resource constraints of edge devices of the IoT, AI/ML-powered blockchain computing is the optimal solution for security. These devices simply cannot handle the heavy-duty software and gateways designed for desktop computers. Autonomic computing, which uses AI and ML, might further improve creative software designs that help with the updating and routine maintenance of IoT devices.
8.2 Edge AI
Recent advancements in AI effectiveness, the proliferation of IoT devices, and the advent of edge computing have unlocked the potential of edge AI. Because of this, edge AI has potential applications in fields like agricultural fertilization, vehicle assistance, and radiology diagnosis that were previously unthinkable A45 . Employing edge AI may improve nearly every industry’s current processes. The fact is that cutting-edge applications are propelling the next wave of AI-driven autonomic computing, which will enhance people’s lives everywhere—at home, in the office, at school, and even while driving. As a term, ”AI at the edge” describes the integration of AI with tangible hardware. ”Edge AI” allows the execution of AI computations near users on the network’s periphery rather than centralizing all an organization’s data in a single location, such as a private data center or a cloud provider’s data center A46 . Automation is a goal for businesses of all sizes because of the many benefits it provides in terms of efficiency, effectiveness, and security. Software that recognizes patterns and has the ability to consistently perform the same actions could prove beneficial in this situation. The world is unpredictable, and human activities encompass endless conditions, making it challenging to completely communicate them in a system of norms and algorithms. Computerization and devices of any kind may now collaborate with the ”intelligence” of human cognition, thanks to advancements in edge AI A47 . It is possible for AI-powered intelligent IoT apps to adapt to new environments and efficiently perform the same or comparable duties. Substantial development in crucial areas has enabled the realistic implementation of AI models at the edge. Advancements in neural networks and other branches of AI have also made universal ML possible. Businesses are increasingly realizing that they can efficiently train AI models and then implement them at the edge A26 . AI at the peripheral requires distributed computing resources. Graphics Processing Units (GPUs), which have recently achieved massive parallelism, power modern AI-driven autonomic systems. The proliferation of IoT-connected gadgets has led to an unprecedented spike in data volume in modern times. The advancement of data collection devices such as sensors, smart cameras, and robotics has enabled the use of AI-driven autonomic computing at the edge in almost every aspect of business A36 . Enhanced reliability, speed, and privacy brought to the battlefield by 5G/6G are equally beneficial to IoT applications.
8.3 Energy Efficiency and Carbon Neutrality
In recent years, there has been a meteoric growth in the amount of data collected and processed. Because of this trend, edge systems have been operating at or near their computational and, consequently, energy consumption limits. Decentralised computing systems like Fog and Edge have emerged as a result of this change A27 . By offloading some of the processing to dispersed edge devices and networks, autonomic cloud computing is able to significantly reduce latency and costs. Apart from the power source, an irregular energy supply presents significant challenges for a wide range of mission-critical and sensing applications. The exponential growth in the variety of IoT devices and the data they generate has challenged the capacity of networking to process, calculate, and transmit data. Similarly, a shortage of computational power, data storage, and electricity currently hinders the development of smaller IoT devices. Therefore, improving the network’s fog and edge nodes’ efficiency is crucial. There has been a recent surge in the importance of data centers’ sustainability and reducing their carbon footprints A49 . Therefore, it is crucial to maintain QoS throughout the process. Despite the setbacks, there have been numerous successes. Researchers employ AI based software, hardware, and transitional methods to address the problem of energy management. Computational frameworks support the development of AI based approaches and procedures that maximize software efficiency. Computing at the edge of a network is one example of this. When it came to hardware, the goal was to minimize power consumption while maximizing performance, especially for this application. Several studies have focused on how to make sensor networks more energy efficient. In the preliminary stage, the researchers employed proactive resource management using AI, scheduling for fog/edge-node hibernation times, and other energy-saving tactics. Regarding the effectiveness and longevity of fog, edge, and cloud platforms, there are numerous unresolved issues and potential avenues for future enhancements A50 . As a result of global warming, the paradigms for autonomic systems that will replace current ones will most likely prioritize electricity and carbon footprints. Systems that assume a constant and steady power supply, a connection to renewable energy sources and alternative methods of reducing energy consumption face more fundamental issues; these problems go beyond the current focus on reducing power consumption per unit of computation. Reduced power usage is necessary to run the CPU and transmitter on energy from the sun or another source. Therefore, reducing the granularity of the fog/edge network could lead to more dispersed, dominant, and robust designs. Studies involving blockchain algorithms and different forms of flexible AI-based autonomic computing could improve energy scheduling in a number of domains, including those dealing with energy limitations.
@@ 9 Conclusions and Summary
The article delves into the relationship between SLA violations and QoS in cloud-based autonomic computing systems, as well as their respective roles. To learn how QoS needs affect self-management, one must be familiar with the development of autonomic cloud computing in order to ascertain the scheduling efficiency of the provided resources. Ultimately, it identified unanswered questions about autonomic cloud computing and other open research topics, which ML and AI systems could potentially address by anticipating resource needs. Importantly, self-adaptive systems driven by AI offer to be environmentally friendly and economically address evolving needs in an unpredictable environment, rather than merely allocating increasing amounts of resources. Consequently, some industry heavyweights are studying and implementing autonomic computing with cutting-edge AI and ML solutions.